Hi. Im studing for a test and im having some problems. I was wondering if anyone can help me. I would really appreciate it. I have 4 questions.
1) Translate into algebraic expression.
The Difference between five times the input and seve.
2) six times the input, added to 8
3) Solve the problem
Write a symbolic rule that represents the relationship in which the output variable t is 5 more than 3.5 times the input variable r. What is the input corresponding to an output value of 26?
4) Write a symbolic rule that represents the relationship in which the output variable t is 7 more than 2.5 times the input variable r. What is the corrisponding to an input value of 8?

Answer by stanbon(75887)

1) Translate into algebraic expression.
The Difference between five times the input and seven.
Let the input be "x":
5x-7
-------------
2) six times the input, added to 8
6x+8
-----------------
3) Solve the problem
Write a symbolic rule that represents the relationship in which the output variable t is 5 more than 3.5 times the input variable r.
t=3.5r+5
----------
What is the input corresponding to an output value of 26?
t=3.5*26 +5
t = 91+5 = 96
-----------------
4) Write a symbolic rule that represents the relationship in which the output variable t is 7 more than 2.5 times the input variable r.
t = 2.5r+7
-----------------
What is the corrisponding to an input value of 8?
t = 2.5*8+7
t = 27
-----------------
Cheers,
Stan H.